OpenStudy (anonymous):

Write two linear equations you can use to graph the equation. y=|3/2x+6|+3

OpenStudy (anonymous):

I will give medal

OpenStudy (lncognlto):

Ok, so, y=|3/2x+6|+3. That means that \[y = \pm(\frac{ 3 }{ 2 }x + 6) + 3\]. Now solving each of these equations, one for the plus and one for the minus, what do we get?

OpenStudy (anonymous):

???

OpenStudy (lncognlto):

The absolute value of 3/2x + 6 is required, so it could be positive or negative.

OpenStudy (anonymous):

kk

OpenStudy (lncognlto):

??

OpenStudy (anonymous):

so it can be negative and positive bcz the absolute valuse

OpenStudy (lncognlto):

I've got to go, but you are in good hands with zzr0ck3r :)

OpenStudy (anonymous):

kk

OpenStudy (zzr0ck3r):

y = |x| implies y=x or y=-x so you have y=|3/2x+6|+3 y-3 = |3/2x+6| so (y-3) = 3/2x+6 and -(y-3) = 3/2x+6 solve for y in each of these equations.

OpenStudy (anonymous):

so slve for y

OpenStudy (anonymous):

*solve

OpenStudy (anonymous):

im lost